OpenStudy (anonymous):

What is the slope of the line which passes through (4, 5) and (0, 1)? undefined 1 0 4

OpenStudy (matt101):

Slope = rise/run = (5 - 1)/(4 - 0) = 4/4 = 1

OpenStudy (lgbasallote):

\[slope = \frac{y_2 - y_1}{x_2 - x_1}\] y2 = 1 y1 = 5 x2 = 0 x1 = 4 therefore... \[slope = \frac{1 - 5}{0-4} = \frac{-4}{-4}\]